OpenStudy (anonymous):

waqas wishes to distribute 128 bananas and also 176 apples equally among a certain number of children. Find the highest number of children who can get the fruits in this way!! ?

OpenStudy (anonymous):

plzzzzzzzzzzzzzzzzzzzzzzzzzzzz help ?

OpenStudy (raden):

i think u can use GCF of 128 and 176

OpenStudy (raden):

hint : 128 = 2^7 and 176=2^4 * 11

OpenStudy (anonymous):

answer in my book is 16 but i dont know how it is 16?

OpenStudy (raden):

yeah, because GCF of 2^7 and 2^4*11 = 2^4

OpenStudy (anonymous):

yeah right!! thankss!!

OpenStudy (raden):

welcome...